# Quantum Down Sampling Filter for Variational Auto-encoder

### v2

@@ Abstract
Variational Auto-encoders (VAEs) are crucial for generative modeling and image reconstruction, but their performance often degrades with low-resolution input (16x16 pixels), leading to blurred or inaccurate results. This study proposes a hybrid model that combines quantum computing in the VAE encoder (Q-VAE) and convolutional neural networks (CNNs) in the decoder. By upscaling the resolution from 16x16 to 32x32 during encoding, we assess the model’s ability to preserve key features while improving image resolution. We tested the model on the MNIST and USPS datasets, comparing it to VAE refered as classical VAE (C-VAE) and the classical direct-passing VAE (CDP-VAE), which uses windowing pooling filters. Evaluation using Fréchet Inception Distance (FID) and Mean Squared Error (MSE) shows that the Q-VAE outperforms both, suggesting its potential for image resolution enhancement and synthetic data generation.
@@ 1 Introduction
Variational Autoencoders (VAEs) have become foundational tools in generative modeling and image reconstruction, especially in tasks that require efficient and precise feature extraction from high-resolution data [1]. VAEs are part of the broader family of latent-variable models [2], wherein an encoder maps input data to a structured latent space, and a decoder reconstructs the data from this representation. Despite their success in a wide range of applications, VAEs often face challenges as latent space is often not expressive enough to model intricate relationships, often leading to blurry or low-quality output [3]. These limitations arise primarily from the difficulty of capturing complex high-resolution dependencies, which become especially apparent when dealing with large data sets (which consists of millions of images) or complex structures [5].
@@
Recent advancements in quantum computing offer promising avenues for overcoming these challenges by exploiting quantum phenomena such as superposition and entanglement. These properties allow for more efficient data representation and processing, potentially overcoming some of the bottlenecks inherent in classical methods for generative modeling [6]. The integration of quantum circuits into neural networks, particularly within the VAE encoder, has shown potential to improve latent space representations, leading to enhanced feature extraction and more accurate reconstructions [7]. Quantum in VAE represents an example of such hybrid models, which leverage quantum operations to better capture the underlying data structure, making them especially well suited for high-resolution image tasks [6].
Recent advances in generative models have led to the increased use of low-resolution images, such as 8x8 or 16x16 pixels, as starting points for training medium-quality output [8]. This approach might seem counterintuitive because of the small size of these images, but it has proven effective for several reasons. Training generative models on low-resolution images allows for faster training times, as smaller images require less computational power and processing, enabling faster iterations and fine-tuning. In addition, low-resolution images force the model to focus on learning high-level patterns and structures rather than fine-grained details. This simplification helps the model build a strong foundation for understanding the essential features of the data. Once the model has learned to generate meaningful, broad structures at low resolution, it can then be used to upscale or refine the output to a medium resolution, such as 32x32, using techniques like super-resolution or progressive refinement. This method allows the model to generate more detailed and sharper images, while avoiding the computational difficulties that come with directly training on medium resolutions like 32x32.
Although 32x32-pixel images are commonly used in many applications, including image classification and generative modeling, classical models like VAEs struggle with these resolutions due to limitations in their architecture [9]. VAEs operate by encoding input data into a compressed probabilistic latent space that is then used to reconstruct the input image. However, with 32x32 images, the encoder has to condense a large amount of pixel information into a smaller latent space, often losing fine details in the process. This leads to blurred or artifact-filled reconstructions, as the latent space might not capture the complexity of the image at this resolution. The encoder’s limited capacity to represent detailed features in the latent space becomes particularly problematic for 32x32 images, which are still complex enough to contain important texture, color variation, and structural information. As a result, VAEs can struggle to generate high-quality images from these lower-resolution inputs [10]. The use of 32x32 resolution is prevalent in many machine learning applications, particularly in image classification tasks and generative modeling, as it strikes a balance between computational efficiency and sufficient detail. Datasets like CIFAR-10 and CIFAR-100, which consist of 32x32 pixel images, are standard benchmarks for evaluating generative models and classifiers [11]. Additionally, modern generative models often employ techniques like progressive training, where low-resolution images (8x8 or 16x16) are initially learned and then progressively refined to medium resolutions, such as 32x32 or beyond. This process stabilizes training and reduces the complexity associated with generating high-resolution outputs [12]. In practice, combining low-resolution training (e.g., 8x8 or 16x16) with upscaling techniques allows generative models to learn more efficiently and handle the complexities of higher-resolution images. By training on smaller images first, the models can focus on understanding the broader patterns in the data, then gradually increase the resolution to produce more detailed output. This method has potential to be useful for reducing computational load and improving the quality of the final images generated.
In this work, we introduce a novel Q-VAE that integrates quantum encoding techniques in the encoder while using Convolutional Neural Networks (CNNs) in the decoder. This hybrid architecture leverages the high expressiveness of quantum states to improve feature extraction and mitigate the blurriness often associated with classical VAEs. Unlike traditional quantum generative models, which typically operate on lower-resolution datasets (e.g. 8x8 or 16x16 pixels) [12], our approach is designed to handle medium-resolution inputs, such as 32x32 pixels, allowing the model to produce better-quality image reconstructions for more complex datasets. In the context of Q-VAE, we use MNIST and USPS datasets primarily due to their simplicity and relevance to early-stage exploration of quantum models in image generation. Both are grayscale datasets, which are less complex than color images, making them ideal for testing new techniques, particularly when working with emerging technologies like quantum computing in machine learning.
@@
Since quantum computing, especially its integration into deep learning models, is still in the early stages, focusing on simpler datasets allows for a more controlled environment where we can better understand the impact of quantum techniques on model performance. Furthermore, the datasets used in this study, such as MNIST and USPS (with image resolutions of 28x28 and 16x16 pixels, respectively), are relatively small in size, which is advantageous when exploring the potential of quantum circuits for feature extraction and latent space representation. Their reduced resolution helps reduce the computational burden compared to higher-resolution images. To ensure a fair evaluation, the MNIST dataset is first down-sampled from 28x28 to 16x16 pixels, and then, like the USPS dataset, it is up-sampled to 32x32 pixels before being used as input to assess our model’s performance.
By focusing on grayscale images, we reduce the complexity of the problem, allowing for clearer insights into how quantum techniques can improve feature extraction and image reconstruction, particularly for simpler tasks like recognizing digits. Once the approach is successfully validated on these datasets, the methods can be extended to more complex and higher-resolution data, such as color images or higher resolutions.
@@
  • Introduces a new model quantum VAE (Q-VAE), where quantum encoding is applied only in the encoder, simplifying the model and improving the feature extraction capabilities.
  • Compares Q-VAE with the classical Direct Passing CDP-VAE, highlighting performance differences due to the inclusion of quantum encoding in the encoder.
@@
  • Investigates whether converting images from low resolution (16x16) to medium resolution (32x32) improves model performance.
The structure of this paper is as follows. Section 2 provides a review of related work on classical VAEs and their integration with quantum computing. Section 3 presents the background on Variational Autoencoders (VAEs). In Section 4, the methodology for the proposed quantum-enhanced VAE model is introduced. Section 5 outlines the numerical simulations conducted, while Section 6 discusses the performance analysis. Finally, Section 7 concludes the paper, summarizing the key findings and suggesting potential avenues for future research.
@@ 2 Related Research
In recent years, variational auto-encoders (VAEs) have garnered significant attention in generative modeling, especially for image reconstruction tasks [21]. Traditional VAEs are praised for their ability to encode data into a structured latent space and regenerate approximations of the input. However, classical VAEs often suffer from blurred reconstructions, [3], prompting research to improve VAE architectures for better image quality and feature extraction [25, 26, 27]. Several studies have focused on enhancing VAE performance using convolutional architectures [43, 30, 31, 32]. Gatopoulos et al. introduced convolutional VAEs to better capture image features, surpassing the performance of fully connected architectures in medium-resolution data sets such as CIFAR-10 and ImageNet32 [28]. Furthermore, combining VAE with adversarial training, such as in Adversarial Autoencoders (AAE) [33], incorporates discriminator networks to enhance the visual quality of generated images, address the blurriness issues inherent in classical VAEs, and improve machine learning applications. In addition, anti-adversarial VAEs [34, 35] and VAE-GAN hybrids [36] enhance image generation by incorporating adversarial training, leading to sharper images by reducing blurriness [37].
Quantum Machine Learning (QML) has emerged as a promising field, with studies exploring the integration of quantum circuits into generative models. Quantum GANs (QGANs) leverage quantum computing to enhance the data representation and image generation capabilities of GANs. Although QGANs are not directly related to VAEs, they represent a quantum-based generative approach that improves data diversity and generation quality [39]. Quantum Variational Autoencoders [40, 41, 42, 43, 44, 45] build on the foundational work by Khoshaman et al., who proposed a quantum VAE where the latent generative process is modeled as a quantum Boltzmann machine (QBM). Their approach demonstrates that quantum VAE can be trained end-to-end using a quantum lower bound to the variational log-likelihood, achieving state-of-the-art performance on the MNIST dataset [38]. They have used Log-Likelihood (LL) and Evidence Lower Bound (ELBO) as the measurement and have shown quantum VAE confidence level for the numerical results is smaller than ±0.2 in all cases. Gircha et al. further advanced the quantum VAE architecture by incorporating quantum circuits in both the encoder, which utilizes multiple rotations, and the decoder [42]. These quantum VAE models primarily focus on leveraging quantum methods in the latent space representation, enhancing the generative capabilities of VAEs.
@@ 3 Background
Variational Auto-encoders (VAEs) are widely recognized as powerful generative models in machine learning. They are based on a probabilistic framework where an encoder network maps input data $\mathbf{x}$ to a latent distribution $q(\mathbf{z}|\mathbf{x})$, and a decoder network reconstructs the data from the latent variables $\mathbf{z}$ as shown in equation 1 [4].
$L_{\theta,\phi}(x)=\mathbb{E}_{q_{\phi}(z|x)}\left[\log p_{\theta}(x,z)-\log q_{\phi}(z|x)\right]$ (1)
where $\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}$ represents the expectation with respect to the approximate posterior $q(\mathbf{z}|\mathbf{x})$. The reconstruction term encourages the model to generate data points that are similar to the observed data, while the KL divergence regularizes the learned latent variables by pushing the approximate posterior towards the prior distribution. Although VAEs have been successful in many domains, including image generation and anomaly detection, they often face significant challenges when applied to high-resolution data. This issue can arise because the latent space is often not expressive enough to model the intricate relationships. [14].
@@ 3.1 Encoder
A classical VAE encoder is typically implemented as a fully connected network or a Convolutional Neural Network (CNN), with the aim of mapping the input data to a lower-resolution latent space. The primary objective of the encoder is to process the input data $x$, transform it through a series of hidden layers, and output two parameters: the mean $\mu$ and the log variance $\log(\sigma^{2})$, which are used for variational inference [1]. In the case of image data, the input image is first reshaped into a vector and passed through multiple fully connected layers. The transformations performed by these layers help extract important features from the input. For example, given an input image of dimension 1024 (flattened from a $32\times 32$ image), the encoder progressively reduces the resolution’s, often mapping the input image $x_{0}$ into a lower-resolution space $x_{1}$.
3.1.1 CNN or Fully Connected Layers in Encoders
CNNs are particularly useful for image-based data because they can capture local spatial patterns (e.g., edges, textures) by using convolution layers. These convolution layers progressively down-sample the input image and extract high-level features, while the fully connected layers act to compress these features into a more compact latent representation. In CNN-based encoders, after the convolution layers extract relevant patterns, fully connected layers are used to map the high-resolution feature maps into the latent space. This approach allows the model to better capture spatial dependencies while still benefiting from the representational power of fully connected layers.
@@ 3.1.2 Windowing and Pooling for Resolution Reduction
Pooling techniques such as max pooling or average pooling play a critical role in reducing the spatial resolution of the input data while retaining important features. In max pooling, for example, the maximum value in a small region (or window) of the input image is selected, effectively reducing the image’s resolution while retaining the most significant features within each region. Average pooling operates similarly, but instead computes the average value of each window. These pooling operations are vital for reducing computational complexity while preserving essential information for later stages of the model. Adaptive pooling is another technique often used in encoders to ensure that feature maps have consistent sizes, regardless of the input size. Adaptive pooling adjusts the pooling operation to produce fixed-sized output maps, which is particularly useful when working with images of varying resolutions. This resolution reduction helps the model focus on more abstract features without losing key information, enabling more efficient learning in the latent space.
@@ 3.2 Latent Space Encoding
Finally, the encoder outputs the mean $\mu$ and the log variance $\log(\sigma^{2})$, which parameterize the variational distribution. These are computed as follows:
@@
where $\mu$ and $\log(\sigma^{2})$ are vectors of dimension latent_dim, which is 16 in our case.
The proposed architecture effectively combines quantum and classical components, leveraging quantum-enhanced feature extraction for downstream Neural Network processing.
@@ 3.3 Decoder
The decoder in a Variational Auto-encoder (VAE) is designed to transform the lower-resolution latent variable $z$, produced by the encoder, back into a high-resolution data space, reconstructing the input data. The decoder typically consists of multiple layers, including fully connected and transposed convolution layers, which progressively reconstruct the input data’s spatial and feature dimensions. The core function of the decoder is to ensure that the reconstructed data match closely the original input, enabling the model to learn both the global structure and the local details of the data. As with the encoder, the decoder is trained to minimize the reconstruction error, often using a loss function that combines both the likelihood of the data given the latent variables and a regularization term that encourages smoothness in the latent-space representation.
@@ 3.3.1 Latent Variable Transformation
The decoding process begins by passing the latent variable $z$ through a fully connected layer to map it to a higher resolution feature space. This transformation creates a tensor that matches the expected dimensions for the next layers in the decoder, such as the transposed convolution layers. The fully connected layer enables the decoder to effectively learn a mapping from the compact latent space back to the data space, ensuring that relevant information is maintained during this transformation.

### v3

@@ Abstract
Variational autoencoders (VAEs) are fundamental for generative modeling and image reconstruction, yet their performance often struggles to maintain high fidelity in reconstructions. This study introduces a hybrid model, quantum variational autoencoder (Q-VAE), which integrates quantum encoding within the encoder while utilizing fully connected layers to extract meaningful representations. The decoder uses transposed convolution layers for up-sampling. The Q-VAE is evaluated against the classical VAE and the classical direct-passing VAE, which utilizes windowed pooling filters. Results on the MNIST and USPS datasets demonstrate that Q-VAE consistently outperforms classical approaches, achieving lower Fréchet inception distance scores, thereby indicating superior image fidelity and enhanced reconstruction quality. These findings highlight the potential of Q-VAE for high-quality synthetic data generation and improved image reconstruction in generative models.
Keywords Quantum Variational Auto-encoder, Image Reconstruction, Convolutional Neural Networks, Resolution Enhancement, Fréchet Inception Distance, Synthetic Data Generation
@@ 1 Introduction
Variational encoders (VAEs) are widely used in generative modeling and image reconstruction, providing an efficient framework to learn meaningful latent representations [1]. VAEs employ an encoder-decoder structure, where the encoder maps input data into a probabilistic latent space, and the decoder reconstructs the input from sampled latent variables. Fidelity in image reconstruction and generation refers to how well the generated images match the original images in terms of structural accuracy [2], detail preservation [3], and perceptual similarity [4]. However, classical VAEs (C-VAEs) often struggle to generate high-fidelity images due to inherent limitations in their latent representation [5]. Additionally, the decoder’s reliance on pixel-wise reconstruction loss leads to blurry and over-smoothed images, restricting their effectiveness in capturing complex data distributions [6].
@@
Recent advancements in quantum computing offer promising avenues for overcoming these challenges by exploiting quantum phenomena such as superposition and entanglement. These advances in quantum computing have introduced promising approaches to enhance deep generative models by leveraging quantum parallelism and superposition [7]. The integration of quantum circuits into neural networks, particularly within the VAE encoder, has shown potential to improve latent space representations, leading to enhanced feature extraction and more accurate reconstructions [8]. Quantum neural networks represent an example of such hybrid models, which leverage quantum operations to better capture the underlying data structure, making them especially well suited for high-dimensional image tasks [9].
@@
This study presents a hybrid model referred to quantum VAE (Q-VAE). Q-VAE combines the quantum down sampling filter, which performs image resolution reduction and quantum encoding, and VAE. This is a novel approach that integrates quantum transformations within the encoder while utilizing fully connected layers for feature extraction. The decoder employs transposed convolution layers for upsampling. Q-VAE is evaluated against C-VAE and classical direct passing VAE (CDP-VAE), which incorporates windowed pooling filters (resolution reduction). Since quantum computing, especially its integration into deep learning models, is still in the early stages, focusing on simpler datasets allows for a more controlled environment where we can better understand the impact of quantum techniques on model performance. Furthermore, the datasets used in this study, such as MNIST [10] and USPS [11] (with image resolutions of 28x28 and 16x16 pixels, respectively), are relatively small in size, which is advantageous when exploring the potential of quantum circuits for feature extraction and latent space representation. Their reduced resolution helps reduce the computational burden compared to higher-resolution images. To ensure a fair evaluation, the MNIST dataset is first down-sampled from 28x28 to 16x16 pixels, and then, like the USPS dataset, it is up-sampled to 32x32 pixels before being used as input to assess our model’s performance. By focusing on grayscale images, we reduce the complexity of the problem, allowing for clearer insights into how quantum techniques can improve feature extraction and image reconstruction, particularly for simpler tasks like recognizing digits. Once the approach is successfully validated on these datasets, the methods can be extended to more complex and higher-resolution data, such as color images or higher resolutions.
@@
  • Introduces a new model, Q-VAE, where quantum encoding is applied only in the encoder, simplifying the model and improving the feature extraction capabilities.
  • Compares Q-VAE with the CDP-VAE, highlighting performance differences due to the inclusion of quantum encoding in the encoder.
@@
  • This study explores the effects of upscaling images to 32×32 on model performance, examining how resolution influences feature extraction and representation.
The structure of this paper is as follows. Section 2 provides a review of related work on classical VAEs and their integration with quantum computing. Section 3 presents the background on VAEs. In Section 4, the methodology for the proposed Q-VAE model is introduced. Section 5 outlines the numerical simulations conducted, while Section 6 discusses the performance analysis. Finally, Section 7 concludes the paper, summarizing the key findings and suggesting potential avenues for future research.
@@ 2 Related Research
In recent years, VAEs have garnered significant attention in generative modeling, especially for image reconstruction tasks [1]. Traditional VAEs are valued for encoding data into a structured latent space and reconstructing input approximations, driving research efforts to enhance their architectures for improved image quality and feature extraction [14, 15, 16]. Several studies have focused on enhancing VAE performance using convolutional architectures [17, 18, 19, 20]. Gatopoulos et al. introduced convolutional VAEs to better capture image features, surpassing the performance of fully connected architectures in medium-resolution data sets such as CIFAR-10 and ImageNet32 [21]. Furthermore, combining VAE with adversarial training, such as in Adversarial Autoencoders (AAE) [22], incorporates discriminator networks to enhance the visual quality of generated images, and improve machine learning applications.
2.1 Quantum Generative Modeling
Quantum Machine Learning (QML) has emerged as a promising field, with studies exploring the integration of quantum circuits into generative models. Quantum GANs (QGANs) leverage quantum computing to enhance the data representation and image generation capabilities of GANs [23, 24, 25, 26]. Quantum VAE [27, 28, 33, 29, 30, 31] on the other hand, is based on the foundational work by Khoshaman et al., who proposed a quantum VAE where the latent generative process is modeled as a quantum Boltzmann machine (QBM). Their approach demonstrates that quantum VAE can be trained end-to-end using a quantum lower bound to the variational log-likelihood, achieving state-of-the-art performance on the MNIST dataset [32]. They have used log-likelihood and Evidence Lower Bound (ELBO) as the measurement and have shown quantum VAE confidence level for the numerical results is smaller than ±0.2 in all cases. Gircha et al. further advanced the quantum VAE architecture by incorporating quantum circuits in both the encoder, which utilizes multiple rotations, and the decoder [33]. These quantum VAE models primarily focus on leveraging quantum methods in the latent space representation, enhancing the generative capabilities of VAEs.
@@ 3 Background
3.1 Variational Auto-encoder
VAEs are powerful generative models in machine learning, built on a probabilistic framework. The encoder network maps input data $x$ to a latent distribution $q_{\phi}(z|x)$, while the decoder network reconstructs the original data from the latent variables $z$. Similar to other variational methods, VAEs are optimized using the ELBO as their objective function. Kingma et al. (2019) introduced a parametric inference model $q_{\phi}(z|x)$, also referred to as an encoder or recognition model. The parameters of this inference model, denoted as $\phi$, are known as the variational parameters. The decoder, with parameters $\theta$, is responsible for generating the data from the latent variables $z$. Thus, the parameters $\phi$ and $\theta$ together define the model, where $\phi$ governs the inference process and $\theta$ governs the generative process.[34].
Due to the non-negativity of the Kullback-Leibler (KL) divergence between $q_{\phi}(z|x)$ and $p_{\theta}(z|x)$, ELBO serves as a lower bound on the logarithmic likelihood of the data. This relationship is expressed as:
$L_{\theta,\phi}(x)=\log p_{\theta}(x)-D_{KL}(q_{\phi}(z|x)||p_{\theta}(z|x))$ (1)
Interestingly, the KL divergence $D_{KL}(q_{\phi}(z|x)||p_{\theta}(z|x))$ determines two key aspects:
  1. 1.
    By definition, it quantifies the divergence of the approximate posterior from the true posterior.
  2. 2.
    It represents the gap between the ELBO $L_{\theta,\phi}(x)$ and the marginal likelihood $\log p_{\theta}(x)$, which is also known as the tightness of the bound.
The closer $q_{\phi}(z|x)$ approximates the true posterior $p_{\theta}(z|x)$ in terms of KL divergence, the smaller this gap becomes, leading to a tighter bound. The reconstruction term encourages the model to generate data points that are similar to the observed data, while the KL divergence regularizes the learned latent variables by pushing the approximate posterior towards the prior distribution. Although VAEs have been successful in many domains, including image generation and anomaly detection, they often face significant challenges when applied to high-resolution data. This issue can arise because the latent space is often not expressive enough to model the intricate relationships. [35].
@@ 3.2 Encoder
A classical VAE encoder is typically implemented as a fully connected network or convolutional neural network (CNN) with the aim of mapping the input data to a lower-resolution latent space. The primary objective of the encoder is to process the input data $x$, transform it through a series of hidden layers, and output two parameters: the mean $\mu$ and the log variance $\log(\sigma^{2})$, which are used for variational inference [36]. In the case of image data, the input image is first reshaped into a vector and passed through multiple fully connected layers. The transformations performed by these layers help extract important features from the input. For example, given an input image of dimension 1024 (flattened from a $32\times 32$ image), the encoder progressively reduces the dimension, often mapping the input image $x_{0}$ into a lower-dimension space $x_{1}$.
3.2.1 Fully Connected Layers in Encoders
Fully connected layers in an encoder are crucial for transforming input data into meaningful representations. They work by connecting every neuron in one layer to every neuron in the next, allowing the model to learn complex relationships between features. In the encoder, these layers typically perform dimensionality reduction, extracting the most relevant information from the input and mapping it to a compact latent space. The use of activation functions like ReLU or Sigmoid introduces non-linearity, enabling the encoder to capture intricate patterns. This learning of complex, non-linear relationships and improving performance in capturing patterns make fully connected layers essential for building efficient and powerful encoders. This approach allows the model to better capture spatial dependencies while still benefiting from the representational power of fully connected layers.
@@ 3.2.2 Windowing and Pooling for Resolution Reduction
Fixed pooling techniques, such as max pooling and average pooling, are particularly effective in encoding processes where spatial reduction is necessary. By consistently applying the same pooling operation across the input, the model can reduce dimensionality, improve efficiency, and retain key features that will be crucial for later processing. In max pooling, the maximum value in a small region (or window) of the input image is selected, reducing the image’s resolution while preserving the most significant features in each region. Average pooling works similarly, but instead computes the average value of each window. These pooling operations help to reduce computational complexity while retaining essential information for later stages of the model.
@@ 3.3 Latent Space Encoding
Finally, the encoder outputs the mean $\mu$ and the log variance $\log(\sigma^{2})$, which parameterize the variational distribution. These values are computed using fully connected layers (fc) as follows:
@@
where $\mu$ and $\log(\sigma^{2})$ are vectors of dimension of the latent space, while $h_{2}$ is a vector or matrix resulting from a series of transformations applied to the input data, and it is used as the input to the fully connected layers that output the parameters of the variational distribution. .
@@ 3.4 Decoder
The decoder in a VAE is designed to transform the lower-resolution latent variable $z$, produced by the encoder, back into a high-dimensional data space, reconstructing the input data. The decoder typically consists of multiple layers, including fully connected and transposed convolution layers, which progressively reconstruct the input data’s spatial and feature dimensions. The core function of the decoder is to ensure that the reconstructed data match closely the original input, enabling the model to learn both the global structure and the local details of the data. As with the encoder, the decoder is trained to minimize the reconstruction error, often using a loss function that combines both the likelihood of the data given the latent variables and a regularization term that encourages smoothness in the latent-space representation.
@@ 3.4.1 Latent Variable Transformation
The decoding process begins by passing the latent variable $z$ through a fully connected layer to map it to a higher-dimensional feature space. This transformation creates a tensor that matches the expected dimensions for the next layers in the decoder, such as the transposed convolution layers. The fully connected layer enables the decoder to effectively learn a mapping from the compact latent space back to the data space, ensuring that relevant information is maintained during this transformation.